OpenStudy (anonymous):

rewrite f(x) = x2 − 8x + 3 in vertex form.

OpenStudy (triciaal):

vertex form f(x) = a(x-h)^2 + k h = -b/(2a)

OpenStudy (triciaal):

standard f(x) = ax^2 + bx + c

OpenStudy (hybrik):

Vertex form is y = a(x - h)^2+ k, when H is X in ordered pair and K is Y in ordered pair. To solve this, we can solve by completing the square. y-3=x2-8x, y-3=(x-4)^2 y-3+16=(x-4)^2 y+13=(x-4)^2 [ y=(x-4)^2-13 ] <-- The vertex form